# Invoice Information Extraction: Methods and Performance Evaluation

### v1

@@ 4.3.1 Overall Performance
In terms of processing efficiency, the Docling method exhibits faster average processing times (10 seconds per document) relative to the LlamaExtractor (30 seconds per document). This suggests that Docling prioritizes lightweight processing, likely due to its simpler OCR and layout-based approach, while the LlamaExtractor’s deep language modeling introduces higher computational overhead but delivers better semantic understanding.
@@
Metric Llama Extractor Docling Method
Overall Performance
Overall Accuracy (%) 94 63
Processing Time (avg/doc) 30 sec 10 sec
Consistency Validation
Consistency Check Pass Rate (%) 93 80
Mathematical Validation Errors 0 20
@@ 4.3.2 Consistency Validation
Furthermore, the LlamaExtractor recorded zero mathematical validation errors, while the documentation exhibited 20 errors, suggesting that the internal reasoning capabilities of the Llama enhance the logical verification of the extracted numeric data, such as totals, taxes and line item calculations.
@@ 4.3.3 Technical Characteristics
Metric Llama Extractor Docling Method
Input Format Raw Text/PDF/JPG Raw Text/PDF/JPG
Model Architecture Large Language Model Layout Analysis + OCR
Pre-processing Required Minimal Required
Computational Resources Medium Low
@@ 4.3.4 Error Analysis
Metric Llama Extractor Docling Method
Primary Error Sources Variable descriptions, numeric formatting Mathematical validation, line item description
Complex Layout Handling Good Fair
Multi-page Document Support Yes Limited
@@ 4.3.5 Practical Considerations
Metric Llama Extractor Docling Method
Setup Complexity Medium Low
Scalability Excellent Fair
Cost Efficiency Medium High
Real-time Processing Yes Yes

### v2

@@ 4.3.1 Overall Performance
In terms of processing efficiency, the Docling method exhibits faster average processing times (10 seconds per page) relative to the LlamaExtractor (30 seconds per page). This suggests that Docling prioritizes lightweight processing, likely due to its simpler OCR and layout-based approach, while the LlamaExtractor’s deep language modeling introduces higher computational overhead but delivers better semantic understanding.
@@
Method Overall Accuracy (%) Processing Time (avg/page) Consistency Check Pass Rate (%) Mathematical Validation Errors(%)
Llama Extractor 94 30 sec 93 5
Docling Method 63 10 sec 80 20
@@ 4.3.2 Consistency Validation
Furthermore, the LlamaExtractor recorded 5% mathematical validation errors, while the Docling exhibited 20% errors, suggesting that the internal reasoning capabilities of the Llama enhance the logical verification of the extracted numeric data, such as totals, taxes and line item calculations.
@@ 4.3.3 Technical Characteristics
Method Input Format Model Architecture Pre-processing Required Computational Resources
Llama Extractor Raw Text/PDF/JPG Large Language Model Minimal Medium
Docling Method Raw Text/PDF/JPG Layout Analysis + OCR Required Low
@@ 4.3.4 Error Analysis
Method Primary Error Sources Complex Layout Handling Multi-page Document Support
Llama Extractor Variable descriptions, numeric formatting Good Yes
Docling Method Mathematical validation, line item description Fair Limited
@@ 4.3.5 Practical Considerations
Method Setup Complexity Scalability Cost Efficiency Real-time Processing
Llama Extractor Medium Excellent Medium Yes
Docling Method Low Fair High Yes